OpenStudy (anonymous):

Solve for 0<=theta<360

OpenStudy (anonymous):

\[0\le \theta <360\] \[3\sec \theta=-2\sqrt{3}\]

OpenStudy (anonymous):

150

OpenStudy (anonymous):

It can't be only one value.

OpenStudy (anonymous):

We know that \[\theta<360\]

OpenStudy (anonymous):

There are 360 integer solutions.

OpenStudy (anonymous):

360 Integer solutions for \[3\sec \theta=-2\sqrt{3}\]?

OpenStudy (anonymous):

no, for 0<=theta<360

OpenStudy (anonymous):

Well, obviously.